# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a prominent technique with the potential to become a mainstream method for 3D representations. It can effectively transform multi-view images into explicit 3D Gaussian representations through efficient training, and achieve real-time rendering of novel views. This survey aims to analyze existing 3DGS-related works from multiple intersecting perspec- tives, including related tasks, technologies, challenges, and opportunities. The primary objective is to provide newcomers with a rapid understanding of the field and to assist researchers in methodically organizing existing technologies and challenges. Specifically, we delve into the optimization, application, and extension of 3DGS, categorizing them based on their focuses or motivations. Additionally, we summarize and classify nine types of technical modules and corresponding improvements identified in existing works. Based on these analyses, we further examine the common challenges and technologies across various tasks, proposing potential research opportunities.
@@
3D Gaussian Splatting, Rendering, Technologies.
1 Introduction
@@
The advent of Neural Radiance Fields (NeRF) has ignited considerable interest in the pursuit of photorealistic 3D content. Despite substantial recent advancements that have markedly enhanced NeRF’s potential for practical applications, its inherent efficiency challenges have remained unresolved. The introduction of 3D Gaussian Splatting (3DGS) has decisively addressed this bottleneck, enabling high-quality real-time ($\geq$30 fps) novel view synthesis at 1080p resolution. This rapid development has quickly garnered significant attention from researchers and has led to a proliferation of related works.
@@
Owing to the efficiency and controllable explicit representation of 3DGS, its applications extend across a diverse array of fields. These include enhancing immersive environments in Virtual Reality (VR) and Augmented Reality (AR), improved spatial awareness in robotics and autonomous systems, advanced visual effects in film and animation, and urban planning and architecture, etc.
@@
Figure 1: The introduction of this survey, with the parts marked in RED indicating the new content compared to existing reviews.
@@
To assist readers in quickly grasping the progress in 3DGS research, we provide a comprehensive survey of 3DGS and its downstream tasks. This survey systematically compiles the most important and recent literature on the subject, offering detailed classifications and discussions of their focuses and motivations. However, we find that a considerable number of similar techniques are inevitably mentioned across different tasks. Therefore, we further summarize and classify the various technical modules of 3DGS, such as initialization, attribute setting, and regularization, etc. Based on this summary of techniques, we aim to help readers elucidate the connections between different technologies and enhance various components of 3DGS to meet their customized tasks. Additionally, we investigate the interrelationships between various downstream tasks and technologies in 3DGS, systematically delineating four principal challenges to facilitate future research in this domain. Finally, we emphasize the limitations of existing research and propose promising avenues to address core challenges and advance this rapidly evolving field.
Although several studies have summarized the recent advances in 3DGS [1, 2, 3], our goal is to systematically discuss and fine-grainedly classify the related tasks and techniques of 3DGS, and analyze the commonalities and challenges between them, as illustrated in Fig. 1. Specifically, the main contributions of this survey are as follows:
@@
    This survey discusses 3DGS and its various derivative tasks, including the optimization, application, and extension of 3DGS. Unlike existing reviews, we provide a more detailed classification based on focuses or motivations, which enables readers to gain a more comprehensive understanding of tasks and establish research directions.
@@
    More importantly, we comprehensively analyze the enhancements in various techniques within 3DGS as presented in existing literature, offering detailed classifications and in-depth discussions. This enables readers to discern the commonalities among various improved techniques, thereby assisting in their application to customized tasks.
@@
    We have published an open-source project on GitHub that compiles 3DGS-related articles, and will continue to add new works and technologies into this project. https://github.com/qqqqqqy0227/awesome-3DGS. We hope that more researchers can use it to access the latest research information.
Figure 2: Structure of this Survey, including Related Tasks, Technologies, Challenges, and Opportunities.
As depicted in Fig.2, the structure of this survey is organized as follows: Sec.2 describes the background and details of 3D Gaussian Splatting (3DGS), emphasizing its advantages over Neural Implicit Fields and Point-based Rendering. Sec.3 focuses on optimizing 3DGS to address challenges encountered during the reconstruction process. Sec.4 summarizes the applications of 3DGS and discusses its implementation in downstream tasks. Sec.5 provides an overview of the extensions of 3DGS, exploring ways to enhance its original capabilities. Sec.6 consolidates the various techniques for improving 3DGS modules. Sec.7 reviews the interrelationships among different tasks as well as technologies and summarizes core challenges. Finally, Sec.8 outlines promising avenues for future research, addressing existing challenges and technical optimizations and Sec.9 summarizes this survey.
It should be noted that this survey does not provide an overview of datasets; detailed information can be found in previous reviews [3, 4]. Additionally, in the first half of the article, we primarily describe how existing works address the downstream tasks of 3DGS, while the second half focuses on technologies. Although some repeated works are mentioned, the points of focus and content differ.
2 Preliminaries
2.1 Neural Implicit Field
Neural implicit field representations have garnered significant attention in recent research [5, 6]. These methodologies conceptualize 2D or 3D signals to be reconstructed as fields within corresponding Euclidean spaces, utilizing discrete samples to train neural networks that approximate these fields. This approach facilitates the reconstruction, interpolation, and extrapolation of the original discrete samples, enabling applications such as the super-resolution of 2D images and novel view synthesis of 3D scenes. In the specific context of 3D reconstruction and novel view synthesis, Neural Radiance Fields (NeRF) [7] leverage neural networks to model the geometry and appearance of 3D scenes as density fields and radiance fields. NeRF employs volumetric rendering to establish a mapping from the 3D field to 2D images, thus enabling the reconstruction of 3D signals from multiple 2D images and facilitating novel view rendering. Among the current state-of-the-art methods in this domain, Mip-NeRF 360 [8] stands out for achieving superior rendering quality, while Instant-NGP [9] is notable for its exceptional training efficiency.
However, neural implicit field methods heavily rely on the volumetric rendering process to obtain rendered pixels. This process requires sampling tens to hundreds of points along each ray and inputting them into the neural network to produce the final imaging result. Consequently, rendering a single 1080p image necessitates on the order of $10^{8}$ neural network forward passes, which often takes several seconds. Although some works employ explicit, discretized structures to store continuous 3D fields, thereby minimizing the dependency on neural networks and accelerating the query process for field representations [9, 10, 11], the vast number of sampling points still incurs extremely high rendering costs. Such methods based on volumetric rendering incapable of achieving real-time rendering, thereby limiting their applicability in downstream tasks.
2.2 Point-based Rendering
Since continuous 3D fields do not distinguish between occupied and unoccupied spaces within a scene, a significant number of sampling points during the volumetric rendering process are located in unoccupied spaces. These sampling points contribute minimally to the final rendering result, leading to low rendering efficiency. In contrast, discrete point cloud representations only record the truly occupied parts of the 3D scene, offering a more efficient and precise way to represent the scene. Rendering based on point clouds relies on rasterization rather than stochastic sampling, allowing the use of modern GPUs for real-time rendering.
However, existing high-quality differentiable rendering methods based on point clouds typically depend on pre-constructed point clouds or require dense point cloud reconstruction. These methods do not further optimize the point cloud structure during training [12, 13, 14], resulting in rendering quality that is highly dependent on the initial point cloud quality and making the final images prone to artifacts or incorrect appearances.
2.3 3D Gaussian Splatting
@@ 2.2 Point-based Rendering
3D Gaussian Splatting [15] combines the advantages of neural implicit field and point-based rendering methods, achieving the high-fidelity rendering quality of the former while maintaining the real-time rendering capability of the latter, as shown in Fig. 3. Specifically, 3DGS defines points in the point cloud as 3D Gaussian primitives with volumetric density:
@@
Figure 3: Pipeline and Related Technologies of 3D Gaussian Splatting.
During rendering, 3DGS projects 3D Gaussian primitives onto the 2D imaging plane using the EWA splatting method [16] and employs $\alpha$ blending to compute the final pixel color. This rendering process, for each pixel, is analogous to the discrete form of volumetric rendering used in neural fields, enabling 3DGS to construct complex scene appearances and achieve high-quality rendering. To achieve high-frame-rate, high-resolution differentiable rendering, 3DGS utilizes a tile-based rasterizer. This rasterizer first divides the image into $16\times 16$ tiles and assigns an index to each tile. For each Gaussian primitive, the rasterizer determines which tiles the primitive’s projection intersects and generates a key-value pair for each intersecting tile: the key is 64-bit, with the upper 32 bits representing the tile index and the lower 32 bits representing the projection depth of the Gaussian primitive. By constructing these key-value pairs, the rasterizer only needs to perform a global sort on all key-value pairs, eliminating the need for additional sorting of primitives for each pixel. After sorting, the key-value pairs derived from each tile reside in contiguous memory intervals. The rendering process for each tile is then managed by a CUDA thread block, with the number of threads within each block matching the number of pixels in the tile. Each thread is responsible for the $\alpha$ blending process for its corresponding pixel, thus completing the final rendering.
@@
3 Optimization of 3D Gaussian Splatting
3.1 Efficiency
@@
Efficiency is one of the core metrics for evaluating 3D reconstruction. In this section, we describe it from three perspectives: storage, training, and rendering efficiency.
3.1.1 Storage Efficiency
@@
3DGS requires millions of different Gaussian primitives to fit the geometry and appearance in a scene, leading to high storage overhead: a typical reconstruction of an outdoor scene often requires several hundred megabytes to multiple gigabytes of explicit storage space. Given that the geometric and appearance attributes of different Gaussian primitives may be highly similar, storing attributes for each primitive individually can lead to potential redundancy.
@@
Therefore, existing works [17, 18, 19] primarily focus on applying Vector Quantization [20] (VQ) techniques to compress large numbers of Gaussian primitives.
@@
Compact3D [18] applies VQ to compress different attributes into four corresponding codebooks and stores the index of each Gaussian in these codebooks to reduce the storage overhead. After establishing the codebooks, the training gradients are copied and backpropagated to the original non-quantized Gaussian parameters via the codebooks, updating both the quantized and non-quantized parameters, and discarding the non-quantized parameters when the training is done. Additionally, Compact3D employs run-length coding to compress the sorted index values further, thereby enhancing storage efficiency. Similarly, the work by Niedermayr et al.[19] proposes a sensitivity-aware vector quantization technique to construct codebooks based on sensitivity-aware k-means[21] and utilizes the DEFLATE compression algorithm [22] to further compress the trained codebooks. After training, it proposes a quantization-aware fine-tuning strategy to recover information lost due to VQ.
Furthermore, some works [23, 24, 25] aim at developing pruning strategies or compressing SH parameters.
@@
Some works [24, 23] consider both simultaneously. LightGaussian [24] introduces a Gaussian pruning strategy based on the global significance score and a distillation strategy for high-degree spherical harmonic parameters. Similarly, the work by Lee et al.[23] introduces a learnable mask to reduce the number of original Gaussians and a unified hash-grid-based appearance field[9] to compress the color parameters.
Unlike the aforementioned works, Self-Organizing Gaussian [25] does not use traditional nontopological VQ codebooks for compressing large numbers of Gaussians. Instead, it employs the concept of self-organizing mapping to map Gaussian attributes into corresponding 2D grids. The topological relationships in the 2D grids reflect those in the original attribute space, allowing compression algorithms for topologically structured 2D data to be applied to the unordered Gaussian primitives.
@@
Additionally, there are works [26, 27] focused on improving efficient Gaussian representations.
@@
Scaffold-GS [26] designs anchors and additional attributes for efficient representation, which have the capability to convert to 3DGS. Based on this representation, Scaffold-GS proposes a set of strategies for the growth and pruning of anchors on multi-resolution voxel grids.
GES [27] introduces the Generalized Exponential (GEF) Mixture to replace Gaussian representations, which has the capability to efficiently fit arbitrary signals. By designing a Fast Differentiable Rasterizer and Frequency-Modulated Image Loss for GEF, GES is capable of utilizing a smaller number of GEF primitives while maintaining performance.
3.1.2 Training Efficiency
@@
Improving training efficiency is also important for 3DGS. DISTWAR [28] introduces an advanced technique aimed at accelerating atomic operations in raster-based differentiable rendering applications, which typically encounter significant bottlenecks during gradient computation due to the high volume of atomic updates. By leveraging intra-warp locality in atomic updates and addressing the variability in atomic traffic among warps, DISTWAR implements warp-level reduction of threads at the SM sub-cores using registers. Furthermore, it dynamically distributes atomic computations between the SM and L2 atomic units. This software-only implementation employs existing warp-level primitives to minimize the number of atomic operations directed to L2, thereby significantly enhancing throughput.
3.1.3 Rendering Efficiency
Real-time rendering is one of the core advantages of Gaussian-based methods. Some works that improve storage efficiency can simultaneously enhance rendering performance, for example, by reducing the number of Gaussian primitives. Here, we discuss additional works contributing to these advancements.
@@ 3.1.3 Rendering Efficiency
After training the 3DGS, the work by [29] involves pre-identifying and excluding unnecessary Gaussian primitives through offline clustering based on their spatial proximity and potential impact on the final rendered 2D image. Furthermore, this work introduces a specialized hardware architecture designed to support this technique, achieving a speedup of 10.7× compared to a GPU.
GSCore [30] proposes a hardware acceleration unit for optimizing the rendering pipeline of 3DGS in radiance field rendering. Based on the analysis of performance bottlenecks in Gaussian sorting and rasterization, GSCore introduces optimization techniques such as Gaussian shape-aware intersection tests, hierarchical sorting, and subtile skipping. Implementing these techniques within GSCore results in a 15.86× average speedup over mobile GPUs.
3.2 Photorealism
Photorealism is also a topic worth attention [31]. 3DGS is expected to achieve realistic rendering in various scenarios. Some [32, 33] focus on optimizing in its original settings.
To alleviate the dependence on SfM initialization, GaussianPro [32] introduces an innovative paradigm for joint 2D-3D training. Building upon the 3D plane definition and patch matching technology, it proposes a progressive Gaussian propagation strategy, which harnesses the consistency of 3D views and projection relationships to refine the rendered 2D depth and normal maps. Following geometric filtering and selection processes, the optimized depth and normal maps are employed for densification and additional supervision, ultimately achieving precise geometric representation.
FreGS [33] transitions the supervision process into the frequency domain and employs amplitude and phase attributes from the 2D discrete Fourier transform to alleviate over-reconstruction in 3DGS. Based on this idea, FreGS introduces a frequency-domain-guided coarse-to-fine annealing technique to eliminate undesirable artifacts.
The sharp decline in multi-scale rendering performance is also a topic worth attention [34, 35].
The work [34] first analyzes the causes of aliasing in the frequency domain at low resolutions and in far-away rendering, utilizing Multi-Scale Gaussians to address this issue. Then, Pixel Coverage is defined to reflect the Gaussian size compared to the current pixel size. Based on this concept, it identifies small Gaussians and aggregates them into larger Gaussian for multi-scale training and selective rendering.
Mip-splatting [35], similarly addressing the issue from the perspective of the sampling rate, introduces a Gaussian low-pass filter based on Nyquist’s theorem to constrain the frequency of the 3D Gaussians according to the maximal sampling rate across all observed samples. To address aliasing and dilation artifacts, Mip-splatting replaces the traditional 2D dilation filter with a 2D Mip filter, which approximates a 2D box filter. Unlike modifications in the training phase, SA-GS [36] operates solely by a 2D scale-adaptive filter during test time, making it applicable to any pretrained 3DGS.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a prominent technique with the potential to become a mainstream method for 3D representations. It can effectively transform multi-view images into explicit 3D Gaussian through efficient training, and achieve real-time rendering of novel views. This survey aims to analyze existing 3DGS-related works from multiple intersecting perspectives, including related tasks, technologies, challenges, and opportunities. The primary objective is to provide newcomers with a rapid understanding of the field and to assist researchers in methodically organizing existing technologies and challenges. Specifically, we delve into the optimization, application, and extension of 3DGS, categorizing them based on their focuses or motivations. Additionally, we summarize and classify nine types of technical modules and corresponding improvements identified in existing works. Based on these analyses, we further examine the common challenges and technologies across various tasks, proposing potential research opportunities.
@@
3D Representations, Rendering, 3D Gaussian Splatting.
I Introduction
@@
The advent of Neural Radiance Fields (NeRF) [1] has ignited considerable interest in the pursuit of photorealistic 3D content. Despite substantial recent advancements that have markedly enhanced NeRF’s potential for practical applications, its inherent efficiency challenges have remained unresolved. The introduction of 3D Gaussian Splatting (3DGS) has decisively addressed this bottleneck, enabling high-quality real-time ($\geq$30 fps) novel view synthesis at 1080p resolution. This rapid development has quickly attracted researchers and led to a proliferation of related works.
@@
Owing to the efficiency and controllable explicit representation of 3DGS, its applications extend across a diverse array of fields. These include enhancing immersive environments in Virtual Reality (VR) and Augmented Reality (AR), improved spatial awareness in robotics and autonomous systems, and urban planning and architecture, etc.
@@
Figure 1: The introduction of this survey, with the RED parts indicating the new content compared to existing reviews.
@@
To assist readers in quickly grasping the progress in 3DGS research, we provide a comprehensive survey of 3DGS and its derivative works. This survey systematically compiles the most important and recent literature on the subject, offering detailed classifications and discussions of their tasks and techniques. In examining the technological commonalities across numerous 3DGS variants, we present a structured analysis of technical improvements in fundamental components of vanilla 3DGS, such as initialization, attribute configurations and regularization, etc. Based on this summary of techniques, we aim to help readers synthesize the connections among different improved techniques and provide approaches to enhance various components of vanilla 3DGS to meet their customized tasks. Moreover, we conduct a systematic investigation into the relationships between downstream tasks and their enabling technologies in 3DGS, identifying and analyzing four fundamental challenges. Through careful examination of these challenges, we propose promising research directions to advance this rapidly evolving field, providing a roadmap for future innovations in 3DGS.
Although some comprehensive reviews have documented recent advances in 3DGS[2, 3, 4], they focus on categorizing and discussing existing works by downstream tasks, overlooking technical connections between different tasks, which leads to redundant discussions. Our work is distinctive in providing discussions at two levels: Tasks and Techniques. Specifically, we categorize and discuss existing downstream tasks according to their different motivations or focuses, rather than reviewing all works sequentially. More significantly, we present a thorough examination of technical improvements implemented across various modules of the vanilla 3DGS by existing variants, enabling readers to establish clear relationships between different research fields that share similar methodological foundations. Building upon these, we further investigate the their underlying commonalities and delineate core challenges and opportunities, as shown in Fig. 1. Through this approach, we aim to synthesize recent technical breakthroughs in the 3DGS field and direct researchers’ attention to the core unique challenges facing 3DGS. The primary contributions of this survey can be summarized as follows:
@@
    This survey discusses 3DGS and its various derivative tasks, including the optimization, application, and extension of 3DGS. We provide a classification and discussion based on focuses or motivations, which enables readers to gain a more comprehensive understanding of tasks and establish research directions.
@@
    More importantly, we comprehensively analyze the enhancements in various techniques within 3DGS across different works, offering detailed classifications and in-depth discussions. This enables readers to discern the interrelationships among various improved techniques, thereby assisting in their application to customized tasks.
@@
    We have published an open-source project on GitHub that compiles 3DGS-related articles, and will continue to add new works and technologies into this project. https://github.com/qqqqqqy0227/awesome-3DGS. We hope that researchers can use it to access the latest research information.
II Preliminaries
In this section, we primarily introduce 3DGS to the readers. More background, including Neural Implicit Fields [5] and Point-based Rendering [6], is described in Supplementary-A.
@@ II Preliminaries
3DGS [7] combines the advantages of neural implicit field and point-based rendering methods, achieving the high-fidelity rendering quality of the former while maintaining the real-time rendering capability of the latter. Specifically, as shown in Fig. 2, 3DGS defines points in the point cloud as 3D Gaussian primitives with volumetric density:
@@
During rendering, 3DGS projects 3D Gaussian primitives onto the 2D imaging plane using the EWA splatting method [8] and employs $\alpha$ blending to compute the final pixel color. This process is described in detail in the Supplementary-A.
III Optimization of 3D Gaussian Splatting
III-A Efficiency
@@
Efficiency is one of the core metrics for evaluating 3D reconstruction [9]. In this section, we describe it from three perspectives: storage, training, and rendering efficiency.
III-A1 Storage Efficiency
@@
3DGS requires millions of different Gaussian primitives to fit the geometry and appearance in a scene, leading to high storage overhead: a typical reconstruction of an outdoor scene often requires several hundred megabytes to multiple gigabytes of explicit storage space. Given that the geometric and appearance attributes of different Gaussian primitives may be highly similar, storing attributes for each primitive individually can lead to potential redundancy. Some quantitative reconstruction results are reported in Table I.
@@
Existing works [10, 11, 12] primarily focus on applying Vector Quantization [13] (VQ) techniques to compress 3DGS.
@@
Among them, Compact3D [11] applies VQ to compress different attributes into four corresponding codebooks and stores the index of each Gaussian in these codebooks to reduce the storage overhead. After establishing the codebooks, the training gradients are copied and backpropagated to the original non-quantized Gaussian parameters via the codebooks, updating both the quantized and non-quantized parameters, and discarding the non-quantized parameters when the training is done. Despite achieving efficient 3DGS compression, these methods inevitably encounter quantization errors following discretization and remain sensitive to hyperparameter configurations.
Figure 2: Pipeline and Technologies of 3D Gaussian Splatting.
Furthermore, some works [14, 15] aim at developing efficient pruning strategies.
@@
LightGaussian [15] introduces a Gaussian pruning strategy based on the global significance score and a distillation strategy for high-degree spherical harmonic parameters. Similarly, the work by Lee et al.[14] introduces a learnable mask to reduce the number of original Gaussians. Such methods heavily rely on determining which primitives are non-essential.
@@
Furthermore, there are works [16, 17, 18, 19] focused on improving efficient Gaussian representations or attributes.
@@
Scaffold-GS [16] designs anchors and additional attributes for efficient representation, which have the capability to convert to 3DGS. Based on this representation, Scaffold-GS proposes a set of strategies for the growth and pruning of anchors on multi-resolution voxel grids. Despite their widespread implementation, these approaches face inherent compression limitations due to their unstructured characteristics, which HAC [19] later mitigated through the incorporation of structured hash grid.
TABLE I: Comparison of compression on MipNeRF360 [20].
Method PSNR↑ SSIM↑ LPIPS↓ Size (MB)↓
3DGS 27.49 0.813 0.222 744.7
Scaffold-GS [16] 27.50 0.806 0.252 253.9
HAC [19] 27.53 0.807 0.238 15.26
Compact-3DGS [21] 27.08 0.798 0.247 48.80
EAGLES [10] 27.15 0.808 0.238 68.89
LightGaussian [15] 27.00 0.799 0.249 44.54
Gaussian-SLAM [18] 26.01 0.772 0.259 23.90
Compact3d [11] 27.16 0.808 0.228 50.30
III-A2 Training Efficiency
@@
Improving training efficiency is also important for 3DGS. DISTWAR [22] introduces an advanced technique aimed at accelerating atomic operations in raster-based differentiable rendering applications, which typically encounter significant bottlenecks during gradient computation due to the high volume of atomic updates. By leveraging intra-warp locality in atomic updates and addressing the variability in atomic traffic among warps, DISTWAR implements warp-level reduction of threads at the SM sub-cores using registers. These strategies enable DISTWAR to achieve an average 2.44× acceleration in performance.
III-A3 Rendering Efficiency
@@
Real-time rendering is one of the core advantages of Gaussian-based methods. Some works that improve storage efficiency can simultaneously enhance rendering performance, for example, by reducing the number of Gaussian primitives. Several studies [23, 24] have specifically addressed this issue. After training the 3DGS, the work by [23] involves pre-identifying and excluding unnecessary Gaussian primitives through offline clustering based on their spatial proximity and potential impact on the final rendered 2D image. Furthermore, this work introduces a specialized hardware architecture designed to support this technique, achieving a speedup of 10.7× compared to a GPU.
III-B Photorealism
Photorealism is also a topic worth attention [25]. 3DGS is expected to achieve realistic rendering in various scenarios. Some [26, 27, 28] focus on optimizing under vanilla settings. Among them, GaussianPro [26] introduces an innovative paradigm for joint 2D-3D training. Building upon the 3D plane definition and patch matching technology, it proposes a progressive Gaussian propagation strategy, which harnesses the consistency of 3D views and projection relationships to refine the rendered 2D depth and normal maps. Although these methods demonstrate superior capabilities in handling artifacts and 3D inconsistencies compared to 3DGS, further exploration is still needed for complex geometric reconstruction. To further mitigate this issue, the work [29] introduces a scalable and efficient N-dimensional Gaussian Mixture Model for fast, accurate high-dimensional modeling without domain-specific heuristics or sacrificing computational efficiency but is limited by its reliance on dense data and conservative culling.
The sharp decline in multi-scale rendering performance is also a topic worth attention [30, 31, 32, 33]. Among them, Mip-splatting [31], addressing the issue from the perspective of the sampling rate, introduces a Gaussian low-pass filter and 2D Mip filter based on Nyquist’s theorem to constrain the frequency of the Gaussians according to the maximal sampling rate across all observed samples. Then, to address the over-smoothing issue in Mip-splatting, Analytic-Splatting [33] proposes a novel method that analytically approximates the integral of 2D Gaussian signals within pixel window areas through a conditioned logistic function. While it enhances detail fidelity, it comes at the expense of efficiency.